OpenStudy (anonymous):

Put the equations in slope-intercept form and then graph it. 3x + 7y = 2 3x - 4y = 13

OpenStudy (ramirocruzo):

Hey @hkskqgywi ......Do you know the standard line equation...... y = mx + c You just need to compare both and plot the graph PS:Mate you have a way difficult name... :p

OpenStudy (ramirocruzo):

Is it done????

OpenStudy (anonymous):

Sorry, I was working on a different problem of mine. No, it isn't because I have a difficult time of this kind of algebra... I know the standard equation I just don't know how to put those equations into that

OpenStudy (ramirocruzo):

Okay..... 3x + 7y = 2 7y=2-3x y=2/7-3/7x So, now on comparing, c=2/7, m=-3/7 Similarly, do it yourself for second one.... 3x - 4y = 13

OpenStudy (anonymous):

So 3x - 4y = 13 4y = 13 + 3x y = 13/4 + 3/4x??

OpenStudy (ramirocruzo):

Yup...Now compare for the values of c & m

OpenStudy (anonymous):

c=13/4 & m=3/4?? But is that many fractions okay to graph?? because my question says to graph them to find the solution for the system of equations..

OpenStudy (ramirocruzo):

Yup...There no problem......Or wait...Your question if augmented might become......Find the solution of the equations....

OpenStudy (anonymous):

all I have to do is put them in slope intercept form and then graph them to FIND the solutions. Which for system of equations I don't know if they want me to shade a part or not... I hate algebra lol.

OpenStudy (ramirocruzo):

Man....You are making me confused too....Just find some values of x & y for each equations & plot their graph....Point where they intercept is the solution......

OpenStudy (anonymous):

I am a girl by the way. Ill just try to graph them and see where they point to find the solution

OpenStudy (ramirocruzo):

Sorrryy....Am really sorry.....Pardon me......

OpenStudy (anonymous):

Isn't the X supposed to be in the other place? y = mX + b (my teacher always told us it was b).

OpenStudy (ramirocruzo):

Its the same c i.e. the intercept on Y-Axis.....I researched & found out that in british systems its c & in US its b.....

OpenStudy (anonymous):

so in the slope intercept form you demonstrated for me (thank you by the way), in y= 2/7 - 3/7x -3/7x is going to be the x-intercept? and then 2/7 is going to be the y-intercept?

OpenStudy (ramirocruzo):

 Oki...Sis...Take a look at this.....Also your solution is x=3 & y=-1

OpenStudy (anonymous):

I have to graph it on this http://taylormath.pbworks.com/f/1193085651/f-206-15-ex-1.gif

OpenStudy (ramirocruzo):

When you plot it...It should look like this...

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (ramirocruzo):

Anytime...^_^